OpenStudy (swaqqout_kid1):

Shannon wants to buy cupcakes for her friend's birthday party. The bakery she goes to sells cupcakes for $3.75 each. The bakery also sells gift cards for $1.75, and Shannon decides to buy one card. If Shannon buys 8 cupcakes, how much money does she spend at the bakery? $30.00 $31.75 $17.75 $44.00 @Godlovesme

OpenStudy (anonymous):

\[3.75x + 1.75 = ?\] Where X is the number of cupcakes bought

OpenStudy (godlovesme):

3.75c+1.75 = total 3.75(8)+ 1.75= total can u solve from there? @SwaqqOut_KID1